OpenStudy (rissyroo118):

in need of major help!!! Which is an equation of the line containing the points (4, 10) and (6, 11) in standard form? (Points : 2) –x + 2y = 16 2x – y = –2 x + 2y = 24 x – 2y = –16

OpenStudy (rissyroo118):

Which shows the equation of the line containing the point (4, 6) and having a slope of 4 in slope-intercept form? (Points : 2)

OpenStudy (rissyroo118):

a= -5/3x+4 b=4x-10 c=4y=x-28 d=-4x+10